OpenStudy (anonymous):

[7.06] Choose the correct simplification of (8x + 5)(4x − 6). Select one: a. 32x2 + 28x + 30 b. 32x2 + 28x − 30 c. 32x2 − 28x + 30 d. 32x2 − 28x − 30

OpenStudy (hoblos):

(8x + 5)(4x − 6) = 8x(4x) + 8x(-6) +5(4x) + 5(-6) = 32x2 - 48x +20x -30 = 32x2 -28x -30